If A and B are any two non empty sets and is proper subset of B if n (A) is 5 then min possible value of n(A symmetric diff. B is)

Grade:7

1 Answers

kkbisht
5 years ago
As A \subset B => Symmetric Difference of A & B=  ( A \Delta B) =(A-B) U (B-A) =\phi U (B-A)= B-A ( As A\subset B=>A-B= \phi)
Therefore n(A\Delta B)= n(B-A)=1 ( in case n(A)=4 and n(B)=5 as A is proper subset of B
 
kkbisht